# FREE-Merging: Fourier Transform for Efficient Model Merging

### v2

@@ Abstract
With the rapid growth of deep learning, there is an increasing availability of open-source models for various tasks. However, single fine-tuned models often fall short of meeting the diverse needs of users. Model merging has thus emerged as an efficient method to integrate the capabilities of existing models into a unified model. Nevertheless, existing model merging methods face challenging trade-offs between performance and deployment costs, primarily due to task interference. For the first time, we reveal that task interference is evident in the frequency domain of model parameters, yet current efforts only focus on spatial domain solutions, which are largely ineffective in addressing frequency domain interference. To mitigate the impact of frequency domain interference, we propose FR-Merging, an innovative method that effectively filters harmful frequency domain interference on the backbone with minimal computational overhead. Since performance loss is inevitable with cost-free methods, we propose a lightweight task-specific expert module that dynamically compensates for information loss during merging. This proposed framework, FREE-Merging (FR-Merging with experts), strikes a balanced trade-off between training cost, inference latency, storage requirements, and performance. We demonstrate the effectiveness of both FR-Merging and FREE-Merging on multiple tasks across CV, NLP, and Multi-Modal domains and show that they can be flexibly adapted to specific needs.
@@ 1 Introduction
In the current era of deep learning, the pretrain-finetune paradigm has emerged as the standard methodology [37, 9, 40]. However, in numerous edge applications, users often lack powerful computing resources and large datasets, preferring ready-to-use models for specific scenarios. The recent growth of open-source platforms like HuggingFace [43] has made this more feasible, allowing easy access to models for various tasks. Yet, practical demands are often not fully covered by a single existing model and are typically a union of several models [52, 23]. However, running separate models for each subtask increases storage and inference costs, especially with large models. Model merging attempts to combine existing models to build a model capable of addressing all target tasks [21, 33, 48]. This method not only mitigates the training costs and data privacy issues associated with multi-task learning (MTL) [55, 20] but also reduces deployment costs, garnering widespread attention.
@@
For backbone merging, the key lies in eliminating task interference [51]. We identify a compelling phenomenon that task interference is evident in the frequency domain of model parameters. And we present in Sec. 4 that frequency-domain interference is strongly correlated with performance. Notably, existing methods, which focus only on spatial-domain operations, fail to address frequency-domain interference, leading to suboptimal outcomes. Experimental evidence reveals that task interference is severe in the low-frequency region. This occurs because, high-frequency signals represent fine-grained variations, while low-frequency parts capture the global structure [18], which is more likely to contain task-specific information that leads to task interference. Leveraging the inherent redundancy in fine-tuned parameters [53, 27, 47], we propose to directly filter out low-frequency parts with severe task interference. As detailed in Sec. 5.2, this high-pass filtering substantially enhances generalization with minimal performance loss, preventing task interference during merging. Our lightweight yet highly effective approach, FR-Merging, is the first to apply Fourier filtering to neural network parameters, showing its potential for broader applications.
@@ 2 Related Work
Model Merging. Model merging integrates existing models to handle multiple tasks without training [21, 23, 49], but it faces task interference challenges [48, 36]. Simple averaging [44] results in great performance degradation. Methods like Fisher-Merging [33] and RegMean [23] determine merging coefficients using various matrices but incur high computational costs or data requirements, limiting their usage in edge deployment. Task Arithmetic [21] merges task vectors instead of weights, while Ties-Merging [48] resolves parameter conflicts, and AdaMerging [52] automates coefficient selection. However, these methods have requirements on data distribution and exhibit limited applicability. Techniques such as DARE [53] and PCB-Merging [12] analyze parameters to mitigate performance drops but still face task conflicts. EMR-Merging [20] and Twin-Merging [29] store task-specific knowledge but require large storage and overlook backbone optimization. Our method applies high-pass filtering to reduce task interference and lightweight experts to enhance performance, achieving a balance between merging costs and effectiveness across various tasks. For further details, please refer to Appendix C.2.
@@ 5 Methodology
This section presents our FREE-Merging: Sec 5.1 explains the need for a two-part design, Sec 5.2 and Sec 5.3 cover the FR-Merging for the backbone and the lightweight expert module, respectively, followed by the FREE-Merging in Sec. 5.4 and a complexity analysis in Sec 5.5.
@@ 5.3 Lightweight Expert Extraction
To compensate for information lost due to discarding low-frequency signals, we utilize lightweight task experts added during inference. As analyzed in Sec. 5.2, low-frequency signals contain task-specific information, but preserving them requires an inefficient inverse Fourier transform during each inference. Thus, we propose a more efficient way.
@@
Although magnitude-based extraction [48] exists, limited research on rescaling hinders compression. As rescaling is crucial to expert extraction, we propose a rescaling method. Let $M(v_{i},d)$ represent the top-$d$ ($d$ is a percentage) parameters for task vector $v_{i}$, $\mu_{i}$ be the rescale factor, and $e(v_{i})$ be the expert for task $t_{i}$. Then we have:
@@
Method Base Model SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
@@ 6.1 Merging Vision Models
Method Base Model Full PEFT Avg.
@@ 6.4 Fourier Transform Analysis
Table 7: Ablation of FREE-Merging components on 8 visual tasks.
@@ 6.5 Ablation Study
Table 8: Transferability analysis of FR-Merging on 8 visual tasks.
@@
Ablation on Task Expert. We conduct an ablation study on the expert extraction way. Table 7 shows that incorporating our experts and rescaling method results in significant improvements, confirming their effectiveness. To validate the superiority of our method, we evaluate expert-only performance with varying parameter proportions, comparing Top-K selection, Bernoulli selection [53], SVD extraction [29], and our Top-K selection with rescaling. Fig 5(b) shows that our method excels, particularly under high sparsity.
Router Analysis. We analyze the proposed router method, aiming for strong merging performance despite sub-optimal routing. Since router inaccuracies may arise from limited training resources or edge deployment issues, our methods maintain performance in real-world scenarios. Using a random router to simulate inaccuracy, we compare our method with existing router-based approaches. Our method outperforms Twin-Merging by 13.5% and EMR-Merging by 8.5% on average, as shown in Fig 5(c), indicating that our methods construct a reasonable backbone and generalizable experts.
@@ 6.6 Transferability Analysis
Our proposed FR-Merging can be seen as a method that can integrate with others, as it addresses frequency-domain task interference that current methods overlook. As shown in Table 8, combining various methods with FR-Merging leads to performance improvements. Therefore, FR-Merging can be integrated with more approaches to achieve greater benefits.
@@ 6.7 Speed and Storage Space Analysis
Method Training Tokens Training Cost Storage Param. Inference Cost (/1000 items) Acc.
@@
This part analyzes training costs, inference speed, and storage as shown in Table 9. FR-Merging achieves strong performance with significantly reducing storage demands without extra training or inference costs, while FREE-Merging slightly increases overhead for better performance. Both of these two methods enhance deep learning deployment.
@@ 7 Conclusion
In this paper, we propose FR-Merging and FREE-Merging to efficiently merge existing models to create a multi-task model. FR-Merging constructs a backbone without training by high-pass filtering to remove information that dominates task interference. Meanwhile, FREE-Merging employs lightweight experts to offset information loss during merging. Our method balances training, storage, inference costs, and performance, showing high practical value.
@@ Appendix D More Discussions and Future Directions
In this section, we first provide a more detailed discussion of task interference in the frequency domain in D.1. Next, we present the experimental results of FR-Merging on language models in D.2 and D.3. Additionally, we discuss the effectiveness and construction of the expert module in D.4 and D.5, as well as the role of the router in D.6. We also explore the selection of the cutoff frequency in D.7 and conclude with a discussion of future works in D.8.
@@ D.1 Task Interference in Frequency Domain
In this section, we further discuss the task interference presented in Sec. 4 and provide additional visualization results.
@@ D.5 Task Experts Construction Analysis
Table 12: Ablation of scaling for experts on 8 visual tasks.
@@ D.7 Filtering Frequency Hyperparameter Analysis
D.8 Future Works
In this section, we outline directions for our future work. First, we aim to extend our merging methods to fine-tuned models with different architectures, enabling broader applicability. Additionally, we plan to further investigate the differences between PEFT and full fine-tuning and design customized merging methods tailored to PEFT, enabling effective utilization and deployment of existing PEFT results.
@@ E.1 Merging Vision Models
In this section, we present the results of merging using ViT-L/14 as the pre-trained model across the 30 vision datasets mentioned in Sec. 6.1, as shown in Table 14.
From the results, we can see that, consistent with the conclusions in the main text, our FR-Merging outperforms all cost-free methods, efficiently constructing a high-performance merged backbone. With the dynamic addition of experts, FREE-Merging achieves satisfactory performance at a minimal additional cost.
@@
Task Arithmetic 30.81 59.86 91.97 73.06 31.05 89.03 9.34 31.25 74.56 93.61
Ties-Merging 23.21 42.82 92.31 73.22 21.09 89.39 5.30 10.98 72.86 91.88
Breadcrumbs 33.95 54.29 91.83 66.24 32.12 14.30 0.19 32.33 69.04 90.92
@@
RegMean 90.71 89.65 99.10 76.14 71.00 93.60 16.28 74.13 86.62 98.54
AdaMerging 81.22 87.54 97.97 75.23 22.76 91.02 0.42 44.60 89.13 96.91
Twin-Merging 92.10 89.34 98.54 81.43 78.45 94.14 58.12 69.43 92.23 98.85
EMR-Merging 93.40 90.23 98.34 83.54 76.54 95.12 54.34 76.64 94.54 98.65
@@
Task Arithmetic 47.65 37.11 33.24 79.59 80.68 39.66 41.86 11.05 14.73 15.50
Ties-Merging 26.03 27.05 12.84 78.27 34.33 6.17 31.28 5.61 3.71 6.79
Breadcrumbs 40.91 35.29 39.16 83.80 81.94 54.73 27.58 3.25 17.81 15.20
@@
RegMean 42.89 83.42 34.62 83.64 95.26 78.94 49.78 48.67 30.53 34.66
AdaMerging 53.09 76.76 48.34 81.98 95.69 68.91 48.19 18.02 16.68 24.83
Twin-Merging 72.65 80.64 68.78 85.49 96.75 74.72 60.47 57.47 41.25 50.57
EMR-Merging 72.34 82.35 67.35 85.73 96.46 76.65 63.11 60.03 40.48 49.57
@@
Task Arithmetic 87.02 84.62 80.20 37.01 17.41 55.88 36.32 51.14 25.23 6.15
Ties-Merging 76.58 67.22 78.61 40.74 10.54 52.69 19.90 19.13 3.91 1.50
Breadcrumbs 79.25 85.15 75.65 38.51 22.89 47.96 26.07 17.303 28.06 1.40
@@
RegMean 98.10 92.58 82.59 56.96 66.13 58.58 57.11 68.74 65.31 19.79
AdaMerging 99.13 93.38 84.19 59.90 25.70 64.09 48.66 66.55 38.54 7.94
Twin-Merging 100.0 94.24 85.57 65.87 75.49 71.25 63.81 71.67 70.82 25.56
EMR-Merging 100.0 95.45 86.43 65.94 76.46 72.45 65.43 73.84 72.24 28.24
@@
E.2 Merging Language Models
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg
Individual 60.18 94.04 89.22 90.63 91.41 87.20 92.71 79.06 85.55
Task Arithmetic 18.78 85.89 79.90 74.03 83.78 59.08 69.67 62.09 66.65
Ties-Merging 20.48 84.40 81.13 58.19 85.70 64.65 74.81 42.96 64.04
Breadcrumbs 25.23 85.21 78.19 68.45 81.07 64.40 79.83 63.18 68.19
PCB-Merging 23.46 84.59 79.36 68.89 81.98 63.57 76.56 64.47 67.86
RegMean 36.67 90.60 75.74 62.68 83.55 70.02 82.35 58.48 70.01
EMR-Merging 37.82 87.91 82.31 72.90 82.61 78.91 83.24 67.92 74.20
Twin-Merging 52.30 93.23 88.65 68.64 81.99 78.80 88.40 74.37 78.29
@@ E.2 Merging Language Models
Fisher Merging 83.3 83.3 56.7 54.2 58.3 83.1 42.2 94.1 45.9 41.0 42.2 62.2
Task Arithmetic 74.1 83.3 62.8 49.1 49.3 87.5 41.5 95.3 60.8 49.4 50.0 63.9
@@
Ties-Merging 78.0 83.3 67.9 57.6 59.7 81.7 42.8 90.3 66.9 51.3 51.1 66.4
Breadcrumbs 71.9 59.4 53.1 31.2 64.1 78.1 46.9 90.2 50.0 31.2 25.0 54.6
PCB-Merging 85.9 83.3 61.9 57.1 63.9 82.4 42.7 91.2 64.2 47.8 45.9 66.1
@@
RegMean 81.2 58.3 53.8 55.2 53.5 80.9 40.1 92.5 43.3 39.2 40.2 58.0
EMR-Merging 81.8 87.5 66.6 56.1 65.3 82.4 44.7 93.6 65.7 43.8 50.8 67.1
Twin-Merging 81.2 85.6 65.4 57.2 66.3 81.6 44.4 92.9 66.5 42.4 51.2 66.8

### v3

@@ Abstract
With the rapid growth of deep learning, there is an increasing availability of open-source models for various tasks. However, single fine-tuned models often fall short of meeting the diverse needs of users. Model merging has thus emerged as an efficient method to integrate the capabilities of existing models into a unified model. Nevertheless, existing model merging methods face challenging trade-offs between performance and deployment costs, primarily due to task interference. For the first time, we reveal that task interference is evident in the frequency domain of model parameters, yet current efforts only focus on spatial domain solutions, which are largely ineffective in addressing frequency domain interference. To mitigate the impact of frequency domain interference, we propose FR-Merging, an innovative method that effectively filters harmful frequency domain interference on the backbone with minimal computational overhead. Since performance loss is inevitable with cost-free methods, we propose a lightweight task-specific expert module that dynamically compensates for information loss during merging. This proposed framework, FREE-Merging (FR-Merging with experts), strikes a balanced trade-off between training cost, inference latency, storage requirements, and performance. We demonstrate the effectiveness of both FR-Merging and FREE-Merging on multiple tasks across CV, NLP, and Multi-Modal domains and show that they can be flexibly adapted to specific needs. Our code is available at: https://github.com/Zhengsh123/FREE-Merging.
@@ 1 Introduction
In the current era of deep learning, the pretrain-finetune paradigm is a standard procedure [37, 9, 40]. However, edge applications often lack powerful computing resources and data, favoring ready-to-use models for specific scenarios. The recent growth of open-source platforms like HuggingFace [43] has made this feasible. Yet, practical demands are often not fully covered by a single existing model and are typically a union of several models [52, 23]. However, running separate models for each subtask increases storage and inference costs, especially with large models. Model merging attempts to combine existing models to build a model capable of addressing all target tasks [21, 33, 48]. This method reduces training costs, data privacy issues in multi-task learning (MTL) [56, 20] and deployment costs, garnering widespread attention.
@@
For backbone merging, the key lies in eliminating task interference [51]. We identify a compelling phenomenon in which task interference is evident in the frequency domain of model parameters. And we present in Sec. 4 that frequency-domain interference is strongly correlated with performance. Notably, existing methods, which focus only on spatial-domain operations, fail to address frequency-domain interference, leading to suboptimal outcomes. Experimental evidence reveals that task interference is severe in the low-frequency region. This occurs because, high-frequency signals represent fine-grained variations, while low-frequency parts capture the global structure [18], which is more likely to contain task-specific information that leads to task interference. Leveraging the inherent redundancy in fine-tuned parameters [53, 27, 47], we propose to directly filter out low-frequency parts with severe task interference. As detailed in Sec. 5.2, this high-pass filtering substantially enhances generalization with minimal performance loss, preventing task interference during merging. Our lightweight yet highly effective approach, FR-Merging, is the first to apply Fourier filtering to neural network parameters, showing its potential for broader applications.
@@ 2 Related Work
Model Merging. Model merging integrates existing models to handle multiple tasks without training [21, 23, 49], but faces task interference challenges [48, 36]. Simple averaging [44] causes great performance degradation. Methods like Fisher-Merging [33] and RegMean [23] use matrices to determine merging coefficients but are computationally expensive or data-intensive, limiting edge deployment. Task Arithmetic [21] merges task vectors instead of weights, while Ties-Merging [48] resolves parameter conflicts, and AdaMerging [52] automates coefficient selection. However, these approaches have data distribution requirements and limited applicability. Techniques such as DARE [53] and PCB-Merging [12] mitigate performance drops but still face task conflicts. EMR-Merging [20] and Twin-Merging [29] store task-specific knowledge but require large storage and neglect backbone optimization. Our method uses high-pass filtering and lightweight experts to balance merging costs and effectiveness across various tasks. For further details, please refer to Appendix C.2.
@@ 5 Methodology
This section presents FREE-Merging: Sec. 5.1 explains the need for a two-part design, Sec. 5.2 and Sec. 5.3 cover the FR-Merging for the backbone and the lightweight expert module, respectively, followed by the FREE-Merging in Sec. 5.4 and a complexity analysis in Sec. 5.5.
@@ 5.3 Lightweight Expert Extraction
To compensate for information lost due to discarding low-frequency signals, we utilize lightweight experts added during inference. As analyzed in Sec. 5.2, low-frequency signals contain task-specific information, but preserving them requires an inefficient inverse Fourier transform during each inference. Thus, we propose a more efficient way.
@@
[…]
@@
Method
Base
Model
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
@@ 6.1 Merging Vision Models
Method
Base
Model
Full PEFT Avg.
@@ 6.4 Fourier Transform Analysis
Table 7: Ablation of FREE-Merging components on 8 visual tasks.
@@ 6.5 Ablation Study
Table 8: Transferability analysis of FR-Merging on 8 visual tasks.
@@
Ablation on Task Expert. We conduct an ablation study on the expert extraction way. Table 7 shows that incorporating our experts and rescaling method results in significant improvements, confirming their effectiveness. To validate the superiority of our method, we evaluate expert-only performance with varying parameter proportions, comparing Top-K selection, Bernoulli selection [53], SVD extraction [29], and our Top-K selection with rescaling. Fig. 5(b) shows that our method excels, particularly under high sparsity.
Router Analysis. We analyze the proposed router method, aiming for strong merging performance despite sub-optimal routing. Since router inaccuracies may arise from limited training resources or edge deployment issues, our methods maintain performance in real-world scenarios. Using a random router to simulate inaccuracy, we compare our method with existing router-based approaches. Our method outperforms Twin-Merging by 13.5% and EMR-Merging by 8.5% on average, as shown in Fig. 5(c), indicating that our methods construct a reasonable backbone and generalizable experts.
@@ 6.6 Transferability Analysis
Our proposed FR-Merging can integrate with other methods, as it addresses frequency-domain task interference that current methods overlook. As shown in Table 8, combining various methods with FR-Merging leads to performance improvements. Therefore, FR-Merging can be integrated with more approaches to achieve greater benefits.
@@ 6.7 Speed and Storage Space Analysis
Method
Training
Tokens
Training
Cost
Storage
Param.
Inference Cost
(/1000 items)
Acc.
@@
This part analyzes training costs, inference speed, and storage as shown in Table 9. FR-Merging achieves strong performance with significantly reducing storage costs without extra training or inference costs, while FREE-Merging slightly increases overhead for better performance. Both of these two methods enhance deep learning deployment.
@@ 7 Conclusion
In this paper, we propose FR-Merging and FREE-Merging to efficiently build a multi-task model by merging existing ones. FR-Merging constructs a backbone without training by high-pass filtering to remove information that dominates task interference. Meanwhile, FREE-Merging employs lightweight experts to offset information loss during merging. Our method balances training, storage, inference costs, and performance, showing high practical value.
Acknowledgements
This work is supported by National Natural Science Foundation of China (NSFC) (62232005, 62202126); the National Key Research and Development Program of China (2021YFB3300502).
@@ Appendix D More Discussions and Future Directions
In this section, we first provide a more detailed discussion of task interference in the frequency domain in D.1. Next, we present the experimental results of FR-Merging on language models in D.2 and D.3. Additionally, we discuss the effectiveness and construction of the expert module in D.4 and D.5, as well as the role of the router in D.6. We also explore the selection of the cutoff frequency in D.7 and explored the implications of forgetting model-specific capabilities in D.8. Finally, we conclude with a discussion of future work in D.9.
@@ D.1 Task Interference in Frequency Domain
In this section, we further discuss the task interference presented in Sec. 4 and provide visualization results.
@@ D.5 Task Experts Construction Analysis
Table 12: Ablation of scaling for experts on 8 visual tasks.
@@ D.6 Role of Router
Table 13: Ablation study of router training.
Training Data Size 0% 0.1% 0.5% 1% 2% 5%
ViT-B/32 78.1 87.0 88.2 89.7 90.0 90.1
ViT-L/14 88.3 92.1 93.0 93.7 93.9 94.1
Meanwhile, in this section, we discuss the amount of data required for training the router. In Tab. 13, we present the average performance across eight tasks using FREE-Merging with different sizes of training data in the fusion experiments on ViT-B/32 and ViT-B/14. It can be observed that the performance improvement becomes marginal when the training data exceeds 1%; therefore, in practical applications, we use only 1% of the data for training. If resources are even more limited, using only 0.1% of the data can still yield acceptable results.
@@ D.7 Filtering Frequency Hyperparameter Analysis
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg
Individual 60.18 94.04 89.22 90.63 91.41 87.20 92.71 79.06 85.55
Task Arithmetic [21] 18.78 85.89 79.90 74.03 83.78 59.08 69.67 62.09 66.65
Ties-Merging [48] 20.48 84.40 81.13 58.19 85.70 64.65 74.81 42.96 64.04
Breadcrumbs [7] 25.23 85.21 78.19 68.45 81.07 64.40 79.83 63.18 68.19
PCB-Merging [12] 23.46 84.59 79.36 68.89 81.98 63.57 76.56 64.47 67.86
RegMean [23] 36.67 90.60 75.74 62.68 83.55 70.02 82.35 58.48 70.01
EMR-Merging [20] 37.82 87.91 82.31 72.90 82.61 78.91 83.24 67.92 74.20
Twin-Merging [29] 52.30 93.23 88.65 68.64 81.99 78.80 88.40 74.37 78.29
@@
D.8 Fuse to Forget
In this section, we discuss the forgetting of task-specific components caused by the merging process. In the setting explored in this paper, we aim to preserve the capabilities of multiple individual models simultaneously, so that the merged model can address multiple tasks. Therefore, in our setting, it is crucial to minimize the loss of task-specific capabilities during the fusion process. To address this issue, we propose FR-Merging and FREE-Merging.
On the other hand, [54] suggests that such forgetting is not necessarily harmful, for example, in reducing model bias and other undesirable effects. In their setup, the goal is to actively forget certain undesirable or sensitive information through merging. Thus, the desired behavior of model fusion differs between the two settings.
However, what remains consistent is that the merging process can lead to a loss of each model’s specific capabilities. As discussed earlier, this is primarily due to potential conflicts with the knowledge from other models.
D.9 Future Works
In this section, we present the directions for our future work. First, we plan to extend our current merging methods to fine-tuned models with different architectures, thereby enhancing the generalizability and applicability of our approach across a wider range of model types. In addition, we intend to conduct a deeper investigation into the differences between parameter-efficient fine-tuning (PEFT) and full fine-tuning, aiming to understand their respective characteristics and behaviors more thoroughly. Based on these insights, we will develop customized merging strategies specifically tailored to PEFT, enabling more effective integration and deployment of existing PEFT-based models.
@@ E.1 Merging Vision Models
In this section, we comprehensively present the detailed results of model merging using ViT-L/14 as the pre-trained backbone model across all 30 diverse vision datasets introduced and described in Sec. 6.1, as shown in Table 16.
From the results, we can see that, consistent with the conclusions presented in the main text, our proposed FR-Merging method significantly outperforms all existing cost-free model merging techniques. It is able to efficiently construct a high-performance merged backbone, demonstrating strong generalization and robustness across various tasks without incurring any additional training or computational overhead. This makes FR-Merging a highly effective solution for model merging in resource-constrained settings. In addition, when we consider the dynamic addition of expert models through FREE-Merging, the method is able to maintain satisfactory performance levels while introducing only a minimal amount of extra computational cost. This flexibility allows for adaptive model expansion depending on the complexity of the task, making FREE-Merging a practical approach for real-world applications where both performance and efficiency are important considerations.
@@
Task Arithmetic [21] 30.81 59.86 91.97 73.06 31.05 89.03 9.34 31.25 74.56 93.61
Ties-Merging [48] 23.21 42.82 92.31 73.22 21.09 89.39 5.30 10.98 72.86 91.88
Breadcrumbs [7] 33.95 54.29 91.83 66.24 32.12 14.30 0.19 32.33 69.04 90.92
@@
RegMean [23] 90.71 89.65 99.10 76.14 71.00 93.60 16.28 74.13 86.62 98.54
AdaMerging [52] 81.22 87.54 97.97 75.23 22.76 91.02 0.42 44.60 89.13 96.91
Twin-Merging [29] 92.10 89.34 98.54 81.43 78.45 94.14 58.12 69.43 92.23 98.85
EMR-Merging [20] 93.40 90.23 98.34 83.54 76.54 95.12 54.34 76.64 94.54 98.65
@@
Task Arithmetic [21] 47.65 37.11 33.24 79.59 80.68 39.66 41.86 11.05 14.73 15.50
Ties-Merging [48] 26.03 27.05 12.84 78.27 34.33 6.17 31.28 5.61 3.71 6.79
Breadcrumbs [7] 40.91 35.29 39.16 83.80 81.94 54.73 27.58 3.25 17.81 15.20
@@
RegMean [23] 42.89 83.42 34.62 83.64 95.26 78.94 49.78 48.67 30.53 34.66
AdaMerging [52] 53.09 76.76 48.34 81.98 95.69 68.91 48.19 18.02 16.68 24.83
Twin-Merging [29] 72.65 80.64 68.78 85.49 96.75 74.72 60.47 57.47 41.25 50.57
EMR-Merging [20] 72.34 82.35 67.35 85.73 96.46 76.65 63.11 60.03 40.48 49.57
@@
Task Arithmetic [21] 87.02 84.62 80.20 37.01 17.41 55.88 36.32 51.14 25.23 6.15
Ties-Merging [48] 76.58 67.22 78.61 40.74 10.54 52.69 19.90 19.13 3.91 1.50
Breadcrumbs [7] 79.25 85.15 75.65 38.51 22.89 47.96 26.07 17.303 28.06 1.40
@@
RegMean [23] 98.10 92.58 82.59 56.96 66.13 58.58 57.11 68.74 65.31 19.79
AdaMerging [52] 99.13 93.38 84.19 59.90 25.70 64.09 48.66 66.55 38.54 7.94
Twin-Merging [29] 100.0 94.24 85.57 65.87 75.49 71.25 63.81 71.67 70.82 25.56
EMR-Merging [20] 100.0 95.45 86.43 65.94 76.46 72.45 65.43 73.84 72.24 28.24
@@
Fisher Merging [33] 83.3 83.3 56.7 54.2 58.3 83.1 42.2 94.1 45.9 41.0 42.2 62.2
Task Arithmetic [21] 74.1 83.3 62.8 49.1 49.3 87.5 41.5 95.3 60.8 49.4 50.0 63.9
@@
Ties-Merging [48] 78.0 83.3 67.9 57.6 59.7 81.7 42.8 90.3 66.9 51.3 51.1 66.4
Breadcrumbs [7] 71.9 59.4 53.1 31.2 64.1 78.1 46.9 90.2 50.0 31.2 25.0 54.6
PCB-Merging [12] 85.9 83.3 61.9 57.1 63.9 82.4 42.7 91.2 64.2 47.8 45.9 66.1
@@
RegMean [23] 81.2 58.3 53.8 55.2 53.5 80.9 40.1 92.5 43.3 39.2 40.2 58.0
EMR-Merging [20] 81.8 87.5 66.6 56.1 65.3 82.4 44.7 93.6 65.7 43.8 50.8 67.1
Twin-Merging [29] 81.2 85.6 65.4 57.2 66.3 81.6 44.4 92.9 66.5 42.4 51.2 66.8